Adv. Math. Econ. 8, 1-31 (2006)

Existence of financial equilibria in a multi-period stochastic economy

Laura Angeloni^{1} and Bernard Cornet^{2}

1 Dipartmento di Matematica e Informatica,Università debli Studi di Pergia,Via Vanvitelli 1,06123,Perugia,Italy

2 Department of Economics,University of Kansas,Lawrence,Kansas 66045-7585,USA

Received: July 27, 2005 Revised: July 27, 2005

JEL classification: D52, C62

Mathematical Subject Classification (2000): 91B50, 91B54

Abstract. We consider the model of a stochastic financial exchange economy with finitely many periods. Time and uncertainty are represented by a finite event-tree 𝔻 and consumers may have constraints on their portfolios. We provide a general existence result of financial equilibria, which allows to cover several important cases of financial structures in the literature with or without constraints on portfolios.

Key words: Incomplete markets, financial equilibria, constrained portfolios, multi-period model